OpenStudy (anonymous):

lhelp!. lisa has 6/7 lb of paperclips. of the paperclips, 2/3 are pink and 1/9 are blue.. how many paperclips are pink and how many are blue?

OpenStudy (anonymous):

(6/7)(2/3) = (4/7)lb are pink (6/7)(1/9) = (2/21)lb are blue.

OpenStudy (mathstudent55):

Can't be answered bec we don't know how much each paper clip weighs.